OpenStudy (anonymous):

Solve the equation for r: 3r-s/4 = 5t

OpenStudy (amistre64):

multiply thru by 4 and divde thru by 3; then rearrange

OpenStudy (amistre64):

might be easier to just divide thru by 3 to get r all alone; then move the s part away

OpenStudy (amistre64):

(3r - s/4 = 5t) /3 r - s/12 = 5t/3 r = 5t/3 + s/12 ; simplify as wanted